40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 337 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  13   Thu Oct 25 00:01:21 2007 ranaSoftware InstallationCDSGEO DV => LIGO DV
Martin Hewitson of GEO600 fame has modified the cool GEO DV
to work with the LIGO NDS system with some NDS advice from Rolf (who's over in Germany this week).

I've moved it onto the 40m CDS system and installed it on the AdhikariLab computer named 'django'. It worked immediately.

I modified the main .m file to include the 40m's NDS server. When you run it you have to include the path to the NDS
client written by Ben Johnson.

The attached is a screenshot of it working on a Mac; it looks as cool on Linux.

Its installed in /cvs/cds/caltech/apps/ligoDV/. In matlab you navigate to that directory and then
type addpath('/cvs/cds/caltech/apps/linux/UNIX_NDS_Client_beta2/') to add the NDS client.
On the Solaris machines, type type addpath('/cvs/cds/caltech/apps/solaris9/UNIX_NDS_Client_beta2/') instead.

Then type ligoDV to start it up. Then click away and have fun.

In the example I've selected
C1:PEM-BS_ACC_EAST_Z
and plotted its specgram.

Big grin
Attachment 1: Picture_1.png
Picture_1.png
  12   Wed Oct 24 08:58:09 2007 steveOtherPSLlaser headtemp is up
C1:PSL-126MOPA_HTEMP is 19.3C

Half of the chiller's air intake was covered by loose paper
Attachment 1: htempup.jpg
htempup.jpg
  11   Wed Oct 24 01:43:32 2007 Andrey RodionovOtherGeneralPDF-file -> Will report about first results for XARM during Wednesday meeting

Here is the pdf-file with some graphs showing first results for XARM optimization.

We will discuss alltogether during our Wednesday meeting which starts at 2.40PM. Probably it would be necessary to project this pdf-file to the big screen,
so someone should bring laptop and probably connect it to the projector. I do not have a laptop.

See you on that meeting.
Attachment 1: Andrey_October_24.pdf
Andrey_October_24.pdf Andrey_October_24.pdf Andrey_October_24.pdf Andrey_October_24.pdf Andrey_October_24.pdf Andrey_October_24.pdf Andrey_October_24.pdf Andrey_October_24.pdf
  10   Tue Oct 23 11:08:20 2007 steveOtherGeneralbrush fires
There are big brush fires around LA
40 days plot show no effect in the 40m lab
Attachment 1: brushfires.jpg
brushfires.jpg
  9   Tue Oct 23 09:01:00 2007 ranaOtherOMCPZT calibration/ transfer function.
Are you sure that the error signal sweep is not saturated on the top ends? This is usually the downfall
of this calibration method.
  8   Mon Oct 22 19:27:14 2007 pkpOtherOMCPZT calibration/ transfer function.
We measured the PZT transfer function by comparing the PZT response of the circuit with the cavity in the loop, with that of the circuit without the cavity in the loop. Basically measure the transfer function of the whole loop with the laser/PZT and Op-amps in it. Then take another measurement of the transfer function of everything else besides the PZT and from both these functions, we can calculate the PZT response.

The calibration was done by using the error signal response to a triangular wave of volts applied to the PZT. A measurement of the slope of the error signal , which has three zero-crossings as the cavity sweeps through the sidebands, gives us the Volts/Hz response. In order to derive a frequency calibration of the x axis, we assume that the first zero crossing corresponds to the first side band (-29.5 MHz) and the third one corresponds with the other sideband (+29.5 MHz). And then by using the fact that we know the response of the cavity to a constant frequency shift, we can use the Volts/Hz measurement to calculate the Volts/nm calibration. The slope that was calculated was 3.2e-6 V/Hz and using the fact that the cavity is 1 m in length and the frequency is 1064 nm, we get a calibration of 0.9022 V/nm.

Attachment 1: calib.pdf
calib.pdf
Attachment 2: calibpzt2.pdf
calibpzt2.pdf
Attachment 3: all2.pdf
all2.pdf
Attachment 4: noPZT2.pdf
noPZT2.pdf
  7   Mon Oct 22 12:02:59 2007 ajwRoutineGeneralSTACIS as microseismic shaker
In case we ever want to use our Stacis systems as shakers, check this:
link
  6   Sat Oct 20 11:54:13 2007 waldmanOtherOMCOMC and OMC-SUS work
[Rich, Chub, Pinkesh, Chris, Sam]

Friday the 18th was a busy day in OMC land. Both DCPDs were mounted to the glass breadboard and the OMC-SUS structure was rebuilt to the point that an aluminum dummy mass is hanging, unbalanced. The OSEMs have not be put on the table cloth yet, but everything is hanging free. As for the DCPDs, if you recall one beam is 3mm off center from the DCPD tombstone. Fortunately, one DCPD is nearly 3mm offcenter from the case in the right direction, so the errors nearly cancel. The DCPD is too high, so the beam isn't quite centered, but they're close. We'll get photos of the beam positions in someday. Also, the DC gain between the two PDs is, at first glance, different by 15%. DCPD1, the one seen in transmission has 315 mV of signal while DCPD2 has 280 mV. Not sure why, could be because of beam alignment or tolerances in the Preamp or the angle incident on the diode or the QE of the diodes. The glass cans have *not* been removed.
  5   Fri Oct 19 16:11:38 2007 pkpOtherOMCOMC PZT response
Sam and I locked the laser to the OMC cavity and looked at the error signal as a function of the voltage applied to the OMC PZT.
Here are two plots showing the response as a function of frequency from 1 kHz to 100 kHz and another high-res response in the region of 4.5 kHz to 10 kHz.
Attachment 1: fullspec.jpg
fullspec.jpg
Attachment 2: 4.5to10.jpg
4.5to10.jpg
Attachment 3: 4.5to10.pdf
4.5to10.pdf
Attachment 4: fullres.pdf
fullres.pdf
  3   Thu Oct 18 15:03:14 2007 ajwRoutineGeneralthis is only a test

  2   Thu Oct 18 14:52:35 2007 ranaRoutineASCtest
test

X-(:P;(:))
  1   Wed Oct 17 18:46:33 2007 ranaConfigurationGeneraleLog Change
This is the first entry in the new 40m eLog.

Its GWs or bust now! Big grin



[Hnull][/Hnull]
ELOG V3.1.3-